# 2INFN, Laboratori Nazionali del Gran Sasso, I-67100 Assergi, Italy 3Universität Heidelberg, Zentrum für Astronomie (ZAH), Institut für Theoretische Astrophysik, Albert Ueberle Str. 2, 69120, Heidelberg, Germany 4Universität Heidelberg, Interdisziplinäres Zentrum für Wissenschaftliches Rechnen, Heidelberg, Germany

### v1

@@
Black Hole - Neutron Star and Binary Neutron Star Mergers from Population III and II stars
Benedetta Mestichelli [Uncaptioned image] benedetta.mestichelli@gssi.it112233    Michela Mapelli [Uncaptioned image] mapelli@uni-heidelberg.de3344556677    Filippo Santoliquido[Uncaptioned image] 1122    Manuel Arca Sedda[Uncaptioned image] 112288
@@
1188
   Lavinia Paiella 11 8 8    Guglielmo Costa 5577    Giuliano Iorio 99    Matthew Mould [Uncaptioned image] 10101111    Veronika Lipatova [Uncaptioned image] 33    Boyuan Liu [Uncaptioned image] 33    Ralf S. Klessen [Uncaptioned image] 334412121313
Population III (Pop. III) stars are expected to be massive and to undergo minimal mass loss due to their lack of metals, making them ideal progenitors of black holes and neutron stars. Here, we investigate the formation and properties of binary neutron star (BNS) and black hole-neutron star (BHNS) mergers originating from Pop. III stars, and compare them to their metal-enriched Population II (Pop. II) counterparts, focusing on their merger rate densities (MRDs), primary masses and delay times. We find that, despite the high merger efficiency of Pop. III BNSs and BHNSs, their low star formation rate results in a MRD at least one order of magnitude lower than that of Pop. II stars. The MRD of Pop. III BNSs peaks at redshift $z\sim 15$, attaining a value $\mathcal{R}_{\rm BNS}(z\sim 15)\sim 15\,\rm Gpc^{-3}\,yr^{-1}$, while the MRD of Pop. III BHNSs is maximum at $z\sim 13$, reaching a value $\mathcal{R}_{\rm BHNS}(z\sim 13)\sim 2\,\rm Gpc^{-3}\,yr^{-1}$. Finally, we observe that the black hole masses of Pop. III BHNS mergers have a nearly flat distribution with a peak at $\sim 20\,{\rm M}_{\odot}$ and extending up to $\sim 50\,{\rm M}_{\odot}$. Black holes in Pop. II BHNS mergers show instead a peak at $\lesssim 15\,{\rm M}_{\odot}$. We consider these predictions in light of recent gravitational-wave observations in the local Universe, finding that a Pop. III origin is preferred relative to Pop. II for some events.
@@ 1 Introduction
Population III (Pop. III) stars formed from metal-free gas at high redshift (Haiman et al. 1996; Tegmark et al. 1997; Yoshida et al. 2003). No conclusive evidence of Pop. III stars exists so far (Rydberg et al. 2013; Schauer et al. 2022; Larkin et al. 2023; Meena et al. 2023; Trussler et al. 2023). Current models suggest that Pop. III stars have a top-heavier initial mass function (IMF) compared to stellar populations in the local Universe (Stacy & Bromm 2013; Susa et al. 2014; Hirano et al. 2015; Jaacks et al. 2019; Sharda et al. 2020; Liu & Bromm 2020a, b; Wollenberg et al. 2020; Chon et al. 2021; Tanikawa et al. 2021b; Jaura et al. 2022; Prole et al. 2022; Klessen & Glover 2023; Liu et al. 2024a). Additionally, because of their lack of metals, we expect these stars to lose a small amount of mass through stellar winds over their lifetimes, when they are not fast rotators (Heger et al. 2002; Kinugawa et al. 2014, 2016; Hartwig et al. 2016; Belczynski et al. 2017; Liu et al. 2021; Tanikawa et al. 2021a, 2022, 2023; Costa et al. 2023; Santoliquido et al. 2023; Tanikawa 2024).
In contrast, early Population II (Pop. II) stars are enriched by the yields of Pop. III stars, but still sufficiently metal poor (metallicity $Z$ up to a few ×$\times{}10^{-4}$, Choudhury et al. 2018; Frebel et al. 2019) that mass loss via line-driven stellar winds is quenched (Kudritzki & Puls 2000; Vink et al. 2001; Sabhahit et al. 2023).
@@
Compact binary mergers from Pop. III star progenitors could be detected by current gravitational-wave (GW) interferometers, if their delay times are sufficiently long. As a matter of fact, one of the possible scenarios for the massive binary black hole (BH) merger GW190521, forming a remnant BH of $\sim 150\,{\rm M}_{\odot}$, is the collapse of a Pop. III binary. Indeed, the negligible mass loss and top-heavy IMF of Pop. III stars might lead to the formation of a population of massive BH and binary BHs (Abbott et al. 2019a, 2020a, 2020b; Liu & Bromm 2020b; Kinugawa et al. 2021; Tanikawa et al. 2021a). The properties of these massive compact objects have been extensively studied through a variety of numerical simulations both in isolation (Tanikawa et al. 2021b, a, 2022, 2023; Costa et al. 2023; Santoliquido et al. 2023; Tanikawa 2024) and in star clusters (Sakurai et al. 2017; Wang et al. 2022; Mestichelli et al. 2024; Liu et al. 2024b; Reinoso et al. 2025; Wu et al. 2025). Furthermore, binary BH mergers from Pop. III stars are expected to be significant sources for third-generation interferometers, such as Cosmic Explorer (Reitze et al. 2019; Evans et al. 2021, 2023), and the Einstein Telescope (Kalogera et al. 2021; Branchesi et al. 2023; Santoliquido et al. 2023, 2024; Abac et al. 2025), as these instruments will be capable of probing the high-redshift Universe, where Pop. III star formation is expected to be prevalent.
@@
Besides binary BHs, Pop. III stars might also be the progenitors of binary neutron stars (BNSs) and BH-neutron star binaries (BHNSs) detectable with current and next-generation interferometers. By the end of the third observation run, the LIGO–Virgo–KAGRA (LVK) collaboration has observed several BHNS and BNS merger candidates (Abbott et al. 2019b, 2020c, 2021, 2023a, 2024). Furthermore, the merger GW230529, detected at the beginning of the fourth observing run, has primary component mass $\sim{3-6}$ M (Abac et al. 2024), inside the previously claimed mass gap between the maximum neutron star and minimum BH mass (Orosz 2003; Özel et al. 2010; Farr et al. 2011).
While current GW detectors only probe BNS and BHNS mergers in the nearby Universe (up to $z\sim 0.2$, Abbott et al. 2023a), next generation interferometers will be sensitive to neutron star mergers in a much larger portion of the Universe, increasing the likelihood of observing BNS and BHNS mergers from Pop. II and III star progenitors (Branchesi et al. 2023; Gupta et al. 2024; Abac et al. 2025).
@@
Here, we present models of BNSs and BHNSs formed from the evolution of Pop. III and Pop. II binary stars. We study the distribution of their masses and delay times, and their merger rate densities. To encompass the uncertainties on Pop. III stars, we probe different combinations of IMFs and orbital parameters, following the same approach as our previous works on binary BHs (Costa et al. 2023; Santoliquido et al. 2023; Mestichelli et al. 2024). The paper is structured as follows. Section 2 presents the initial conditions and the numerical codes used. Section 3 shows our main results. Section 4 provides a more detailed analysis of the properties of BHNS mergers and compares formation scenarios involving Pop. III and Pop. II progenitors for currently observed LVK events using Bayes factors. Finally, Section 5 summarizes our conclusions.
@@ 2.1 Population synthesis with sevn
We model the formation of BNSs and BHNSs with the binary population synthesis code sevn(Spera et al. 2019; Mapelli et al. 2020; Iorio et al. 2023), which calculates single stellar evolution by interpolating a set of pre-computed single stellar tracks and describes the main binary interaction processes (mass transfer via stellar winds, Roche lobe overflow, common envelope, tides, and GW decay) by means of semi-analytic prescriptions (Hurley et al. 2002; Iorio et al. 2023). For the common-envelope phase, we set the efficiency parameter $\alpha_{\rm CE}=1$ and model the envelope binding energy $\lambda_{\rm CE}$ as done by Claeys et al. (2014). We refer to Costa et al. (2023), Santoliquido et al. (2023) and Mestichelli et al. (2024) for further details about our models.
2.2 Compact objects formation
According to the prescriptions in sevn , depending on the final carbon-oxygen (CO) core mass ($M_{\rm CO,f}$), a star can form a white dwarf ($M_{\rm CO,f}<1.38\,\rm M_{\odot}$), undergo electron-capture supernova and form a neutron star ($1.38\leq M_{\rm CO,f}<1.44\,\rm M_{\odot}$; Giacobbo & Mapelli 2019), or a core-collapse supernova and form either a neutron star or a BH (CCSN; $M_{\rm CO,f}\geq 1.44\,\rm M_{\odot}$). For CCSNe, we adopt the rapid formalism by Fryer et al. (2012), according to which if $6\leq M_{\rm CO,f}<7\,\rm M_{\odot}$ or $M_{\rm CO,f}\geq 11\,\rm M_{\odot}$, the star will collapse directly into a BH. As a consequence of the adopted CCSN model, neutron star masses fall between $1.19$ and $2\,{\rm M}_{\odot}$, and a lower mass gap between $2$ and $5\,{\rm M}_{\odot}$ naturally emerges. Given the lack of a predictive model for the mass function of neutron stars able to match the masses of Galactic neutron stars (Sgalletta et al. 2023), we do not include neutron star masses in our analysis, for both Pop. III and Pop. II progenitors.
For (pulsational) pair-instability supernovae, we used the fitting formulas reported by Mapelli et al. (2020) and based on the simulations by Woosley (2017). If the pre-supernova He-core mass of a star, $M_{\rm He,f}$, is between 32 and 64 ${\rm M}_{\odot}$, it undergoes pulsational pair-instability. A pair-instability supernova is triggered for $64\leq M_{\rm He,f}\leq 135\,{\rm M}_{\odot}$. Above $135\,{\rm M}_{\odot}$, the star collapses directly to a BH as a consequence of photodisintegration (Woosley 2017; Spera & Mapelli 2017; Renzo et al. 2020).
@@ 2.3 Supernova kicks
After a supernova (SN), compact objects receive a natal kick, that we model following Giacobbo & Mapelli (2020)
@@
with $\langle M_{\rm NS}\rangle$ and $\langle M_{\rm ej}\rangle$ computed from a population of isolated neutron stars with metallicities representative of the Milky Way, $M_{\rm rem}$ the mass of the compact object, and $M_{\rm ej}$ the ejected mass. $f_{\rm H05}$ is a number randomly drawn from a Maxwellian distribution with one-dimensional root mean square $\sigma_{\rm kick}=265\,\mathrm{km\,s^{-1}}$, derived from the proper motions of young Galactic pulsars (Hobbs et al. 2005). This empirical formalism, while approximated, allows us to match the observations of young Galactic pulsars (Hobbs et al. 2005; Verbunt et al. 2017), and results in lower kicks for BHs, in accordance with observations (Atri et al. 2019). In particular, it implies that BHs forming through direct collapse receive a null natal kick.
In a binary system, natal kicks affect the orbital properties, relative orbital velocity, and center of mass of the binary (Hurley et al. 2002). After the kick, the orbital parameters are updated taking into account the new relative orbital velocity and total mass of the binary. For further details, we refer to Iorio et al. (2023).
@@ 2.4 Stellar tracks
Our Pop. III star tracks do not include rotation and are calculated with the stellar-evolution code parsec (Bressan et al. 2012; Chen et al. 2015; Costa et al. 2019, 2025). At the beginning of their main sequence, Pop. III stars cannot ignite the carbon-nitrogen-oxygen (CNO) tri-cycle because of the initial lack of these elements. In order to contrast the gravitational collapse with the energy provided by the proton-proton (pp) chain, Pop. III stars need to reach very high central temperatures. The temperature can become so high that some carbon is synthesized via triple – $\alpha$ reaction (He burning), even during the main sequence. As a consequence, the CNO tri-cycle ignites and replaces the pp chain as the main source of energy of the star (Marigo et al. 2001; Murphy et al. 2021). At the end of the main sequence, Pop. III stars have high enough central temperatures to transition smoothly to the core-He burning phase. Usually, these stellar evolution features appear at $Z<10^{-10}$ (Cassisi & Castellani 1993). As a consequence, for Pop. III stars we assume an initial hydrogen abundance $X=0.751$, He abundance $Y=0.2485$ (Komatsu et al. 2011), and metallicity $Z=10^{-11}$ (Tanikawa et al. 2021b; Costa et al. 2023). For Pop. II stars we adopt a metallicity $Z=10^{-4}$.
We consider stars with zero-age main sequence (ZAMS) mass $m_{\rm ZAMS}$ between 2 and 600 ${\rm M}_{\odot}$. Stars with $2\leq m_{\rm{ZAMS}}<10\,\rm M_{\odot}$ reach the early-asymptotic giant branch phase, while stars with $m_{\rm{ZAMS}}\geq 10\,\rm M_{\odot}$ evolve until late phases of core-oxygen burning, or the beginning of the pair-instability regime. All tracks are computed with the same setup as Costa et al. (2021), including stellar winds, nuclear reaction network, opacity, and equation of state. Above the convective core, we assume a penetrative overshooting with a characteristic parameter of $\lambda_{\rm ov}=0.5$ in units of pressure scale height. For further information on the evolutionary tracks, we refer to Costa et al. (2023).
In Figure 1 we show the evolution of the radii of Pop. III and Pop. II stars. This Figure highlights the compactness of metal-free stars compared to Pop. II stars, especially in the mass range $8\leq m_{\rm ZAMS}\leq 30\,{\rm M}_{\odot}$. As already pointed out by Costa et al. (2023), Pop. III stars with $m_{\rm ZAMS}<100\,{\rm M}_{\odot}$ begin burning He in their cores before their Pop. II counterparts. This difference becomes less evident at higher masses. We also notice that Pop. III stars with $m_{\rm ZAMS}>100\,{\rm M}_{\odot}$ reach the pre-SN stage as blue supergiant stars, while Pop. II stars explode as very large red supergiant stars.

### v2

@@
Black hole–neutron star and binary neutron star mergers from Population III and II stars
Benedetta Mestichelli [Uncaptioned image] benedetta.mestichelli@gssi.it    Michela Mapelli [Uncaptioned image] mapelli@uni-heidelberg.de    Filippo Santoliquido[Uncaptioned image]    Manuel Arca Sedda[Uncaptioned image]
@@
   Lavinia Paiella    Guglielmo Costa    Giuliano Iorio    Matthew Mould [Uncaptioned image]    Veronika Lipatova [Uncaptioned image]    Boyuan Liu [Uncaptioned image]    Ralf S. Klessen [Uncaptioned image]
Population III (Pop. III) stars are expected to be massive and to undergo minimal mass loss due to their lack of metals, making them ideal progenitors of black holes and neutron stars. Here, we investigate the formation and properties of binary neutron star (BNS) and black hole–neutron star (BHNS) mergers originating from Pop. III stars, and compare them to their metal-enriched Population II (Pop. II) counterparts, focusing on their merger rate densities (MRDs), primary masses, and delay times. We find that, despite the high merger efficiency of Pop. III BNSs and BHNSs, their low star formation rate results in a MRD at least one order of magnitude lower than that of Pop. II stars. The MRD of Pop. III BNSs peaks at redshift $z\sim 15$, attaining a value $\mathcal{R}_{\rm BNS}(z\sim 15)\sim 15\,\rm Gpc^{-3}\,yr^{-1}$, while the MRD of Pop. III BHNSs is maximum at $z\sim 13$, reaching a value of $\mathcal{R}_{\rm BHNS}(z\sim 13)\sim 2\,\rm Gpc^{-3}\,yr^{-1}$. Finally, we observe that the black hole masses of Pop. III BHNS mergers have a nearly flat distribution, with a peak at $\sim 20\,{\rm M}_{\odot}$ and extending up to $\sim 50\,{\rm M}_{\odot}$. Black holes in Pop. II BHNS mergers instead show a peak at $\lesssim 15\,{\rm M}_{\odot}$. We consider these predictions in light of recent gravitational-wave observations in the local Universe, finding that a Pop. III origin is preferred relative to Pop. II for some events.
@@ 1 Introduction
Population III (Pop. III) stars formed from metal-free gas at high redshift (haiman1996; tegmark1997; yoshida2003). No conclusive evidence of Pop. III stars exists so far (rydberg2013; schauer2022; larkin2023; meena2023; trussler2023). Current models suggest that Pop. III stars have a top-heavier initial mass function (IMF) compared to stellar populations in the local Universe (sb2013; susa2014; hirano2015; jaacks2019; sharda2020; liu2020; liu2020b; wollenberg2020; chon2021; tanikawa2021; jaura2022; prole2022; klessen2023; liu2024b). Additionally, because of their lack of metals, we expect these stars to lose a small amount of mass through stellar winds over their lifetimes, when they are not fast rotators (heger2002; kinugawa2014; kinugawa2016; hartwig2016; belczynski2017; liu2021b; tanikawa2021b; tanikawa2022; tanikawa2023; costa2023; santoliquido2023; tanikawa2024). In contrast, early Population II (Pop. II) stars are enriched by the yields of Pop. III stars, but still sufficiently metal-poor (with metallicities, $Z$, of up to a few $\times{}10^{-4}$, smc2018; frebel2019) that mass loss via line-driven stellar winds is quenched (kudritzki2000; vink2001; sabhahit2023).
@@
Compact binary mergers from Pop. III star progenitors could be detected by current gravitational-wave (GW) interferometers, if their delay times are sufficiently long. As a matter of fact, one of the possible scenarios for the massive binary black hole (BH) merger GW190521, forming a remnant BH of $\sim 150\,{\rm M}_{\odot}$, is the collapse of a Pop. III binary. Indeed, the negligible mass loss and top-heavy IMF of Pop. III stars might lead to the formation of a population of massive BHs and binary BHs (gw190521; abbott2020a; abbott2020b; liu2020b; kinugawa2021; tanikawa2021b). The properties of these massive compact objects have been extensively studied through a variety of numerical simulations, both in isolation (tanikawa2021; tanikawa2021b; tanikawa2022; tanikawa2023; costa2023; santoliquido2023; tanikawa2024) and in star clusters (sakurai2017; wang2022; mestichelli2024; liu2024; shuai2024; reinoso2025; wu2025). Furthermore, binary BH mergers from Pop. III stars are expected to be significant sources for third-generation interferometers, such as Cosmic Explorer (ce_2019; evans2021; evans2023) and the Einstein Telescope (kalogera2021; branchesi2023; santoliquido2023; santoliquido2024; etblue2025), as these instruments will be capable of probing the high-redshift Universe, where Pop. III star formation is expected to be prevalent.
@@
Besides binary BHs, Pop. III stars might also be the progenitors of binary neutron stars (BNSs) and BH–neutron star binaries (BHNSs) detectable with current and next-generation interferometers. By the end of the third observation run, the LIGO–Virgo–KAGRA (LVK) collaboration had observed several BHNS and BNS merger candidates (gwtc1; gw190814; gw200105_gw200115; gwtc3_2023; gwtc2.1). Furthermore, the merger GW230529, detected at the beginning of the fourth observing run, has a primary component mass of $\sim{3-6}$ M (gw230529), inside the previously claimed mass gap between the maximum neutron star and minimum BH mass (orosz2003; ozel2010; farr2011). While current GW detectors only probe BNS and BHNS mergers in the nearby Universe (up to $z\sim 0.2$, gwtc3_2023), next-generation interferometers will be sensitive to neutron star mergers in a much larger portion of the Universe, increasing the likelihood of observing BNS and BHNS mergers from Pop. II and III star progenitors (branchesi2023; gupta2024; etblue2025).
@@
Here, we present models of BNSs and BHNSs formed from the evolution of Pop. III and Pop. II binary stars. We study the distribution of their masses and delay times, and their merger rate densities. To encompass the uncertainties on Pop. III stars, we probe different combinations of IMFs and orbital parameters, following the same approach as our previous works on binary BHs (costa2023; santoliquido2023; mestichelli2024). The paper is structured as follows. Section 2 presents the initial conditions and the numerical codes used. Section 3 shows our main results. Section 4 provides a more detailed analysis of the properties of BHNS mergers and compares formation scenarios involving Pop. III and Pop. II progenitors for currently observed LVK events using Bayes factors. Finally, Section 5 summarizes our conclusions.
@@ 2.1 Population synthesis with sevn
We modeled the formation of BNSs and BHNSs with the binary population synthesis code sevn(spera2019; mapelli2020; sevn2023), which calculates single stellar evolution by interpolating a set of precomputed single stellar tracks and describes the main binary interaction processes (mass transfer via stellar winds, Roche lobe overflow, common envelope, tides, and GW decay) by means of semi-analytic prescriptions (hurley2002; sevn2023). For the common-envelope phase, we set the efficiency parameter, $\alpha_{\rm CE}=1$, and model the envelope binding energy, $\lambda_{\rm CE}$, as was done by claeys2014. We refer to costa2023, santoliquido2023, and mestichelli2024 for further details about our models.
2.2 Compact object formation
According to the prescriptions in sevn , depending on the final carbon-oxygen (CO) core mass ($M_{\rm CO,f}$), a star can form a white dwarf ($M_{\rm CO,f}<1.38\,\rm M_{\odot}$), undergo electron-capture supernova (SN), and form a neutron star ($1.38\leq M_{\rm CO,f}<1.44\,\rm M_{\odot}$; giacobbo2019), or undergo core-collapse SN and form either a neutron star or a BH (CCSN; $M_{\rm CO,f}\geq 1.44\,\rm M_{\odot}$). For CCSNe, we adopted the rapid formalism by fryer2012, according to which if $6\leq M_{\rm CO,f}<7\,\rm M_{\odot}$ or $M_{\rm CO,f}\geq 11\,\rm M_{\odot}$, the star will collapse directly into a BH. As a consequence of the adopted CCSN model, neutron star masses fall between $1.19$ and $2\,{\rm M}_{\odot}$, and a lower mass gap between $2$ and $5\,{\rm M}_{\odot}$ naturally emerges. Given the lack of a predictive model for the mass function of neutron stars able to match the masses of Galactic neutron stars (vigna_gomez2018; sgalletta2023), we do not include neutron star masses in our analysis, for both Pop. III and Pop. II progenitors.
For (pulsational) pair-instability SNe, we used the fitting formulas reported by mapelli2020 and based on the simulations by woosley2017. If the pre-SN He-core mass of a star, $M_{\rm He,f}$, is between 32 and 64 $\rm M$$_{\odot}$, it undergoes pulsational pair-instability. A pair-instability SN is triggered for $64\leq M_{\rm He,f}\leq 135\,{\rm M}_{\odot}$. Above $135\,{\rm M}_{\odot}$, the star collapses directly to a BH as a consequence of photodisintegration (woosley2017; spera2017; renzo2020).
@@ 2.3 Supernova kicks
After a SN, compact objects receive a natal kick, which we modeled following giacobbo2020:
@@
with $\langle M_{\rm NS}\rangle$ and $\langle M_{\rm ej}\rangle$ computed from a population of isolated neutron stars with metallicities representative of the Milky Way, $M_{\rm rem}$ the mass of the compact object, and $M_{\rm ej}$ the ejected mass. $f_{\rm H05}$ is a number randomly drawn from a Maxwellian distribution with the one-dimensional root mean square $\sigma_{\rm kick}=265\,\mathrm{km\,s^{-1}}$, derived from the proper motions of young Galactic pulsars (hobbs2005). This empirical formalism, while approximated, allows us to match the observations of young Galactic pulsars (hobbs2005; verbunt2017), and results in lower kicks for BHs, in accordance with observations (atri2019). In particular, it implies that BHs forming through direct collapse receive a null natal kick.
In a binary system, natal kicks affect the orbital properties, relative orbital velocity, and center of mass of the binary (hurley2002). After the kick, the orbital parameters are updated, taking into account the new relative orbital velocity and total mass of the binary. For further details, we refer to sevn2023.
@@ 2.4 Stellar tracks
Our Pop. III star tracks do not include rotation and were calculated with the stellar-evolution code parsec (bressan2012; chen2015; costa2019; costa2025). At the beginning of their main sequence, Pop. III stars cannot ignite the carbon-nitrogen-oxygen (CNO) tri-cycle because of the initial lack of these elements. In order to contrast the gravitational collapse with the energy provided by the proton-proton (pp) chain, Pop. III stars need to reach very high central temperatures. The temperature can become so high that some carbon is synthesized via a triple-$\alpha$ reaction (He burning), even during the main sequence. As a consequence, the CNO tri-cycle ignites and replaces the pp chain as the main source of energy of the star (marigo2001; murphy2021). At the end of the main sequence, Pop. III stars have high enough central temperatures to transition smoothly to the core-He burning phase. Usually, these stellar evolution features appear at $Z<10^{-10}$ (cassisi1993). As a consequence, for Pop. III stars we assumed an initial hydrogen abundance of $X=0.751$, He abundance of $Y=0.2485$ (komatsu2011), and metallicity of $Z=10^{-11}$ (tanikawa2021; costa2023). For Pop. II stars we adopted a metallicity $Z=10^{-4}$.
We considered stars with a zero-age main sequence (ZAMS) mass, $m_{\rm ZAMS}$, of between 2 and 600 $\rm M$$_{\odot}$. Stars with $2\leq m_{\rm{ZAMS}}<10\,\rm M_{\odot}$ reach the early-asymptotic giant branch phase, while stars with $m_{\rm{ZAMS}}\geq 10\,\rm M_{\odot}$ evolve until late phases of core-oxygen burning, or the beginning of the pair-instability regime. All tracks were computed with the same setup as in costa2021, including stellar winds, a nuclear reaction network, opacity, and the equation of state. Above the convective core, we assumed a penetrative overshooting with a characteristic parameter of $\lambda_{\rm ov}=0.5$ in units of pressure scale height. For further information on the evolutionary tracks, we refer to costa2023.
In Figure 1 we show the evolution of the radii of Pop. III and Pop. II stars. This figure highlights the compactness of metal-free stars compared to Pop. II stars, especially in the mass range $8\leq m_{\rm ZAMS}\leq 30\,{\rm M}_{\odot}$. As has already been pointed out by costa2023, Pop. III stars with $m_{\rm ZAMS}<100\,{\rm M}_{\odot}$ begin burning He in their cores before their Pop. II counterparts. This difference becomes less evident at higher masses. We also notice that Pop. III stars with $m_{\rm ZAMS}>100\,{\rm M}_{\odot}$ reach the pre-SN stage as blue supergiant stars, while Pop. II stars explode as very large red supergiant stars.